### v1

# SO(3) real algebra method for finite baryon-number density QCD

Abstract
For lattice QCD calculations at finite baryon-number densities, we propose the “SO(3) real algebra method”, in which the SU(3) gauge variable is divided into the SO(3) and SU(3)/SO(3) parts. In this method, we introduce the “maximal SO(3) gauge” by minimizing the SU(3)/SO(3) part of the SU(3) gauge variable. In the Monte Carlo calculation, the SO(3) real algebra method employs the SO(3) fermionic determinant, i.e., the fermionic determinant of the SO(3) part of the SU(3) gauge variable, in the maximal SO(3) gauge, as well as the positive SU(3) gauge action factor $e^{-S_{G}}$. Here, the SO(3) fermionic determinant is real, and it is non-negative for the even-number flavor case ($N_{f}=2n$) of the same quark mass, e.g., $m_{u}=m_{d}$. The SO(3) real algebra method alternates between the maximal SO(3) gauge fixing and Monte Carlo updates on the SO(3) determinant and $e^{-S_{G}}$. After the most importance sampling, the ratio of the SU(3) and SO(3) fermionic determinants is treated as a weight factor. If the phase factor of the ratio does not fluctuate significantly in the sampled gauge configurations, then lattice QCD calculations at finite densities would be feasible.
pacs:
11.15.Ha, 12.12.38.Gc, 12.38.Mh
I Introduction
Quantum chromodynamics (QCD) has been established as the fundamental theory of the strong interaction since the field-theoretical proof of asymptotic freedom in QCD in 1973 [1, 2, 3], In fact, perturbative QCD [4] successfully describes high-energy hadron reactions within the framework of the parton model [5]. However, solving QCD in the low-energy region is quite difficult due to its strong-coupling nature, which leads to nonperturbative properties such as color confinement [6] and spontaneous chiral-symmetry breaking [7].
The lattice QCD formalism [8, 9] was developed as a robust nonperturbative method based on Euclidean QCD, and M. Creutz first performed lattice QCD Monte Carlo calculations around 1980 [10, 11]. The fermion degrees of freedom can be included using the hybrid Monte Carlo algorithm [12]. Finite temperature QCD can be also studied in lattice QCD by imposing temporal (anti)periodicity [13]. Nowadays, lattice QCD is a powerful and reliable method for analyzing nonperturbative QCD. Numerous lattice QCD studies have clarified various aspects of nonperturbative QCD [13] and finite temperature QCD including the QCD transition [14].
Nevertheless, some aspects of nonperturbative QCD remain poorly understood due to the limitations of lattice QCD. One such example is SU(3) color QCD at finite baryon densities, because lattice QCD Monte Carlo calculations break down due to the sign problem [15].
To begin with, we briefly review SU(3) Euclidean QCD and the current status of high density QCD. QCD is an SU(3) gauge theory described by the quark field $\psi(x)$ and the gluon field,
$A_{\mu}(x)=\sum_{a=1}^{8}A_{\mu}^{a}(x)T^{a}\in{\rm su}(3),\quad A_{\mu}^{a}(x)\in{\bf R},$ (1)
where $T^{a}$ (a=1,2,,8a=1,2,...,8) denote the SU(3) generators on the color degrees of freedom. In the Euclidean metric, the QCD action $S[A,\psi,\bar{\psi}]$ is written as
$S[A,\psi,\bar{\psi}]=S_{G}[A]+\int d^{4}x~\bar{\psi}(\not\!\!D+m)\psi\in{\bf R},$ (2)
with the SU(3) covariant derivative $D_{\mu}\equiv\partial_{\mu}+igA_{\mu}$. (In this paper, we use hermite $\gamma$-matrices of $\gamma_{\mu}^{\dagger}=\gamma_{\mu}$ in the Euclidean metric.) The gauge action $S_{G}[A]$ only contains the gluon field strength $G_{\mu\nu}=G_{\mu\nu}^{a}T^{a}\in{\rm su}(3)$ with $G_{\mu\nu}^{a}\in{\bf R}$, and it is real and non-negative in the Euclidain metric:
$S_{G}[A]\equiv\int d^{4}x~\frac{1}{2}{\rm Tr}(G_{\mu\nu}G_{\mu\nu})=\int d^{4}x~\frac{1}{4}G_{\mu\nu}^{a}G_{\mu\nu}^{a}$ (3)
$\geq 0.$ (4)
At zero density, the Euclidean QCD generating functional is expressed by
$Z$ $=$ $\int DAD\psi D\bar{\psi}e^{-S[A,\psi,\bar{\psi}]}$ (5)
$=$ $\int DAe^{-S_{G}[A]}{\rm Det}(\not\!\!D+m),$ (6)
where the gauge action factor is positive, $e^{-S[A]}>0$, and the (zero-density) fermionic determinant is real,
${\rm Det}(\not\!\!D+m)\in{\bf R}.$ (7)
In fact, one can derive
$[{\rm Det}(\not\!\!D+m)]^{*}={\rm Det}[(\not\!\!D+m)^{\dagger}]={\rm Det}(-\not\!\!D+m)$ (8)
$=$ ${\rm Det}[\gamma_{5}(-\not\!\!D+m)\gamma_{5}]={\rm Det}(\not\!\!D+m),$ (9)
using the anti-hermitian nature of the covariant derivative, $\not\!\!D^{\dagger}=-\not\!\!D$, and the anticommutable nature of $\gamma_{\mu}$ and $\gamma_{5}$. Then, for the even-number flavor case of the same quark mass, the fermionic determinant is non-negative. In this case, the gauge action factor and the fermionic determinant can be identified as probability densities in the Monte Carlo method, and lattice QCD calculations can be performed at zero density. Lots of studies of the QCD vacuum and hadrons have been conducted using lattice QCD Monte Carlo calculations for both zero-temperature [13] and finite-temperature systems [14].
Finite baryon-number density QCD is also important for understanding the QCD phase diagram [16, 17], nuclear systems, and neutron stars based on the fundamental principles of the strong interaction. However, a serious problem called the sign problem prevents practical calculations of SU(3) lattice QCD at finite densities [15]. In lattice QCD, many methods have been proposed to overcome the sign problem and to clarify the QCD phase diagram with a quark chemical potential $\mu$ and temperature $T$.
For instance, the Taylor expansion method [18, 19] employs a Taylor expansion with respect to $\mu/T$ and is effective in regions of small chemical potentials, such as $\mu<T$. The complex Langevin method [20, 21, 22, 23] is one of the hopeful methods because its statistical sampling methods can handle complex interactions without encountering algorithmic difficulties and enable nonperturbative calculations, although the theoretical foundation for the convergence is incomplete and the correctness of solutions must be verified for each dynamical system [24, 25]. The reweighing method [26, 27] is a useful general approach to deal with controllable and difficult factors separately [28] and has been applied to finite density QCD [29]. An analytic continuation has been attempted from systems with pure imaginary chemical potentials $\mu\in i{\bf R}$ [30, 31, 32], where no sign problem appears. The Lefschetz thimble method [33] and its applications have been investigated as an interesting approach to high density QCD. This method uses a complex path where the imaginary part of the action is constant or optimized to control the sign problem [34, 35, 36, 37, 38]. For heavy quark systems, a sign-problem-free effective model constructed by combining strong coupling and hopping parameter expansion has been investigated as an indirect approach to finite density QCD [39]. Despite the development of calculation techniques mentioned above, no conclusive method has yet been established.
Instead, many authors have studied systems that resemble finite density QCD without the sign problem. For instance, since SU(2) is pseudo-real, SU(2) color QCD does not encounter the sign problem, and high-density systems have been studied in lattice QCD [41, 40, 42, 43, 44], although baryons are bosons in SU(2) QCD. Even in SU(3) QCD, no sign problem appears for the systems with isospin chemical potentials [45, 46] for the $N_{f}=2$ case with the same u, d quark mass, and lattice QCD simulations have been performed [47]. This type of system could be found in the mismatch of the proton-to-neutron ratio, as is the case inside neutron stars. Also, in SU(3) QCD, no sign problem appears for systems with a chiral chemical potential [48] as a result of the chiral magnetic effect [49]. Such a system may be realized in the presence of strong magnetic fields [50, 51, 52]. Although these studies may provide insight into the qualitative behavior of high density QCD, these examples are rather special or differ significantly from real finite-density SU(3) color QCD.
In this paper, we propose a new method of the “SO(3) real algebra method” for lattice QCD calculations at finite baryon-number density. This paper is organized as follows. Section II provides a brief review of finite density QCD and the sign problem, a significant difficulty associated with Monte Carlo calculations. In Sec. III, we investigate the SO(3) subalgebra/subgroup of the SU(3) algebra/group, as well as the SU(3)/SO(3) coset space. In Sec. IV, we present factorization of an SU(3) group element into an SO(3) element and an SU(3)/SO(3) element, which results in an SO(3) projection. In Sec. V, we introduce a new gauge of the “maximal SO(3) gauge” in SU(3) QCD. In Sec. VI, we formulate the maximal SO(3) gauge in SU(3) lattice QCD. In Sec. VII, we present the formalism of the “SO(3) real algebra method” for continuum QCD at finite baryon-number density. In Sec. VIII, we formulate the SO(3) real algebra method in SU(3) lattice QCD. Section IX is devoted to a summary and conclusion.
II Finite baryon-number density QCD
In this section, we briefly review finite density SU(3) QCD with a quark-number chemical potential $\mu$, which is equal to one third of the baryon-number chemical potential. In the physical case, the chemical potential $\mu$ is real, $\mu\in{\bf R}$. However, here it is treated as a general complex number, $\mu\in{\bf C}$.
The origin of the sign problem lies in the complex value of the action or the fermionic determinant for SU(3) QCD at finite densities, even in the Euclidean metric. In the Euclidean metric, the finite-density SU(3) QCD action $S[A,\psi,\bar{\psi};\mu]$ with the quark-number chemical potential $\mu$ is generally complex,
$S[A,\psi,\bar{\psi};\mu]=S_{G}[A]+\int d^{4}x~\bar{\psi}(\not\!\!D+m+\mu\gamma_{4})\psi$ (10)
$\in{\bf C}.$ (11)
Accordingly, the SU(3) fermionic determinant $\Delta_{\rm FD}[A](\mu)$ is generally complex in finite density QCD,
$\Delta_{\rm FD}[A](\mu)\equiv{\rm Det}(\not\!\!D+m+\mu\gamma_{4})\in{\bf C},$ (12)
after integrating out the fermion field in the QCD generating functional,
$Z(\mu)$ $=$ $\int DAD\psi D\bar{\psi}e^{-S[A,\psi,\bar{\psi};\mu]}$ (13)
$=$ $\int DAe^{-S_{G}[A]}{\rm Det}(\not\!\!D+m+\mu\gamma_{4}).$ (14)
Therefore, in the case of finite densities, unlike ordinary lattice QCD calculations, it is impossible to identify the action factor $e^{-S[A,\psi,\bar{\psi};\mu]}\in{\bf C}$ or the fermionic determinant $\Delta_{\rm FD}[A](\mu)\in{\bf C}$ as a probability density in the Monte Carlo calculation.
In finite density SU(3) QCD, although the generating functional $Z$ itself is real, the phase factor of the fermionic determinant fluctuates violently and cancels itself out, making numerical simulations impossible. This is the sign problem.
However, if the phase factor is sufficiently suppressed during the most importance sampling, Monte Carlo simulations could be performed. Note again that the generating functional $Z$ itself is real, and therefore the sampled gauge configurations can, in principle, be chosen so that their determinant is close to being real.
To examine the complex conjugate of the fermionic determinant $\Delta_{\rm FD}[A](\mu)$ in finite density QCD, we use the Dirac representation for the Euclidean (hermite) $\gamma$-matrices, in which only $\gamma_{2}$ is purely complex, $\gamma_{2}^{*}=-\gamma_{2}$. In the Dirac representation, the complex conjugation matrix $C$ is expressed as
$C\equiv i\gamma_{5}\gamma_{2},$ (15)
and it satisfies
$\gamma_{\mu}^{*}=C\gamma_{\mu}C^{-1}\quad(\mu=1,2,3,4).$ (16)
The complex conjugate of the SU(3) fermionic determinant $\Delta_{\rm FD}[A](\mu)$ can be written as
$\Delta_{\rm FD}[A](\mu)^{*}$ $=$ ${\rm Det}\{\gamma_{\alpha}(\partial_{\alpha}+igA_{\alpha})+m+\mu\gamma_{4}\}^{*}$ (17)
$=$ ${\rm Det}\{\gamma_{\alpha}^{*}(\partial_{\alpha}-igA_{\alpha}^{*})+m+\mu^{*}\gamma_{4}^{*}\}$ (18)
$=$ ${\rm Det}[C\{\gamma_{\alpha}(\partial_{\alpha}-igA_{\alpha}^{*})+m+\mu^{*}\gamma_{4}\}C^{-1}]$ (19)
$=$ ${\rm Det}\{\gamma_{\alpha}(\partial_{\alpha}-igA_{\alpha}^{*})+m+\mu^{*}\gamma_{4}\}$ (20)
$=$ $\Delta_{\rm FD}[-A^{*}](\mu^{*}).$ (21)
In fact, there is an important relation of
$\Delta_{\rm FD}[A](\mu)^{*}=\Delta_{\rm FD}[-A^{*}](\mu^{*}),$ (22)
and it is generally complex in SU(3) QCD for the physical case of $\mu\in{\bf R}$.
In this paper, we focus on the property of
$-A_{\rm SO(3)}^{\mu}(x)^{*}=A_{\rm SO(3)}^{\mu}(x)$ (23)
for the SO(3) gauge field
$A_{\rm SO(3)}^{\mu}(x)\equiv\sum_{i=2,5,7}A^{\mu}_{i}(x)T^{i}\in{\rm so}(3)\subset{\rm su}(3),$ (24)
which is extracted from the SU(3) gauge field $A_{\mu}(x)=\sum_{a=1}^{8}A_{\mu}^{a}(x)T^{a}\in{\rm su}(3)$ in QCD. In fact, as far as the SO(3) gauge part $A_{\rm SO(3)}$ is concerned, the fermionic determinant is real as
$\Delta_{\rm FD}[A_{\rm SO(3)}](\mu)^{*}$ $=$ $\Delta_{\rm FD}[-A_{\rm SO(3)}^{*}](\mu^{*})$ (25)
$=$ $\Delta_{\rm FD}[A_{\rm SO(3)}](\mu)\ \in{\bf R}$ (26)
for the physical case of real finite densities, $\mu\in{\bf R}$.
In this paper, we explore the possibility of suppressing the net phase factor of the SU(3) fermionic determinant of the sampled configurations by using a new gauge, the “maximal SO(3) gauge”, which maximally forces the sampled SU(3) gauge configurations to closely resemble real ones.
Toward lattice QCD calculations at finite densities, we propose a new method, the “SO(3) real algebra method”, using the SO(3) part of SU(3) QCD in the maximal SO(3) gauge.
III SO(3) subalgebra/subgroup in SU(3) and SU(3)/SO(3) coset space
For our new concepts of the “maximal SO(3) gauge” and the “SO(3) real algebra method”, properties of SO(3) and SU(3)/SO(3) play an essential role. In this section, after reviewing the SO(3) subalgebra of the SU(3) algebra, we investigate the SO(3) subgroup of the SU(3) group, as well as the SU(3)/SO(3) coset space.
III.1 SO(3) subalgebra of SU(3) algebra
We use the SU(3) algebra in the fundamental representation, where the SU(3) generators $T^{a}$ are expressed using the Gell-Mann matrices $\lambda^{a}$,
$T^{a}=\frac{\lambda^{a}}{2}\quad(a=1,2,3,...,8).$ (27)
The SU(3) generators $T^{a}$ satisfy the algebra relation
$[T^{a},T^{b}]=if^{abc}T^{c}$ (28)
with the structure constant $f^{abc}\in{\bf R}$.
The SU(3) algebra contains an SO(3) subalgebra. We divide the SU(3) generators $T^{a}(a=1,2,3,...,8)$ into the SO(3) generators $T^{i}(i=2,5,7)$ and the SU(3)/SO(3) generators $T^{\bar{i}}({\bar{i}}=1,3,4,6,8)$.
In this paper, we denote the SO(3) generators by $T^{i}~(i\equiv 2,5,7)$, which are explicitly defined by
$T^{2}$ $=$ $\frac{1}{2}\begin{pmatrix}0&-i&0\\ i&0&0\\ 0&0&0\end{pmatrix},\quad T^{5}=\frac{1}{2}\begin{pmatrix}0&0&-i\\ 0&0&0\\ i&0&0\end{pmatrix}$, (29)
$T^{7}$ $=$ $\frac{1}{2}\begin{pmatrix}0&0&0\\ 0&0&-i\\ 0&i&0\end{pmatrix}$. (30)
On the other hand, we define the SU(3)/SO(3) generators as $T^{\bar{i}}~({\bar{i}}\equiv 1,3,4,6,8)$, where the indices are barred, and they are explicitly defined by
$T^{1}$ $=$ $\frac{1}{2}\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&0\end{pmatrix},\quad T^{3}=\frac{1}{2}\begin{pmatrix}1&0&0\\ 0&-1&0\\ 0&0&0\end{pmatrix}$, (31)
$T^{4}$ $=$ $\frac{1}{2}\begin{pmatrix}0&0&1\\ 0&0&0\\ 1&0&0\end{pmatrix},\quad T^{6}=\frac{1}{2}\begin{pmatrix}0&0&0\\ 0&0&1\\ 0&1&0\end{pmatrix}$, (32)
$T^{8}$ $=$ $\frac{1}{2\sqrt{3}}\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&-2\end{pmatrix}$. (33)
For simplicity, we fix the above-mentioned representation in this paper.
In this representation, the SO(3) generators $T^{i}$ satisfy
$T^{i*}=~^{t}T^{i}=-T^{i}\quad(i=2,5,7),$ (34)
and the SU(3)/SO(3) generators $T^{\bar{i}}$ satisfy
$T^{\bar{i}*}=~^{t}T^{\bar{i}}=T^{\bar{i}}\quad(\bar{i}=1,3,4,6,8).$ (35)
Next, in terms of SO(3) and SU(3)/SO(3) generators, we consider the SU(3) structure constant $f^{abc}\in{\bf R}$ (a,b,c=1,2,8a,b,c=1,2...,8) satisfying the SU(3) algebra relation (28). The SU(3) structure constant $f^{abc}$ is a perfect anti-symmetric tensor, and its components are expressed as
$f_{123}$ $=$ $2,$ (36)
$f_{147}$ $=$ $-f_{156}=f_{246}=f_{257}=f_{345}=-f_{367}=1,$ (37)
$f_{458}$ $=$ $f_{678}=\sqrt{3}.$ (38)
Regarding the structure constant related to the SO(3) algebra, we note the useful relation
$f^{\bar{i}ij}=-2i{\rm Tr}(T^{\bar{i}}[T^{i},T^{j}])$ $=$ $0$ (39)
$({\bar{i}}=1,3,4,6,8$ ; $i,j=2,5,7)$ (40)
because $T^{i},T^{j}]$ is an SO(3) algebra. From the structure constant (38), we also find
$f^{\bar{i}\bar{j}\bar{k}}=0\quad({\bar{i}},{\bar{j}},{\bar{k}}=1,3,4,6,8)$ (41)
for the SU(3)/SO(3) generators.
In this paper, we define the “SO(3) index” as $i\equiv 2,5,7$, and we use the symbol $i$ to represent it. Similarly, we define the “SU(3)/SO(3) index” as ${\bar{i}}\equiv 1,3,4,6,8$, and we use the anti-index symbol $\bar{i}$} to represent it. Additionally, repeated indices of $i$ or $\bar{i}$} of mean the summation over SO(3) or SU(3)/SO(3) indices, respectively.
III.2 SO(3) subgroup and SU(3)/SO(3) coset space in SU(3) group
In this subsection, we investigate the SO(3) subgroup and the SU(3)/SO(3) coset space in the SU(3) group.
Using three real parameters $\theta^{i}$, the SO(3) group element $u$ is expressed as
$u=e^{i\theta^{i}T^{i}}\in{\rm SO(3)},\quad\theta^{i}\in{\bf R}~~(i=2,5,7).$ (42)
As an important property, the SO(3) group element $u$ is real,
$u^{*}=u,\quad{\rm i.e.,}~~u^{\dagger}=~^{t}u,$ (43)
because of $(i\theta^{i}T^{i})^{*}=i\theta^{i}T^{i}.$
The representative element of the SU(3)/SO(3) group element $M$ is chosen to be
$M=e^{i\theta^{\bar{i}}T^{\bar{i}}}$ $\in$ ${\rm SU(3)/SO(3)},$ (45)
$\theta^{\bar{i}}\in{\bf R}~~(\bar{i}=1,3,4,6,8)$
using five real parameters $\theta^{\bar{i}}$. The SU(3)/SO(3) group element $M$ is a symmetric matrix,
$^{t}M=M,\quad{\rm i.e.,}~M^{\dagger}=M^{*},$ (46)
because of $^{t}(\theta^{\bar{i}}T^{\bar{i}})=\theta^{\bar{i}}T^{\bar{i}}$ and $^{t}[(\theta^{\bar{i}}T^{\bar{i}})^{n}]=(\theta^{\bar{i}}T^{\bar{i}})^{n}$ for $n=0,1,2,$...
For any symmetric matrix $\cal S$} and any anti-symmetric matrix $\cal A$}, the trace of their product is zero, ${\rm Tr}({\cal SA})=0$:
${\rm Tr}({\cal SA})={\cal S}_{ij}{\cal A}_{ji}={\cal S}_{ji}(-{\cal A}_{ij})=-{\rm Tr}({\cal SA})=0.$ (47)
Therefore, the SU(3)/SO(3) group element $M$ satisfies
${\rm Tr}(MT^{i})=0\quad(i=2,5,7)$ (48)
because $M$ is symmetric and the SO(3) generators $T^{i}$ ($i=2,5,7$) are anti-symmetric.
III.3 SO(3) subgroup element
For the SO(3) subalgebra, we also use the alternative form of the generators $\Lambda^{a}$ ($a=1,2,3$) defined by
$\Lambda^{1}$ $\equiv$ $\lambda^{7}=\begin{pmatrix}0&0&0\\ 0&0&-i\\ 0&i&0\end{pmatrix},\quad\Lambda^{2}\equiv-\lambda^{5}=\begin{pmatrix}0&0&i\\ 0&0&0\\ -i&0&0\end{pmatrix}$, (49)
$\Lambda^{3}$ $\equiv$ $\lambda^{2}=\begin{pmatrix}0&-i&0\\ i&0&0\\ 0&0&0\end{pmatrix}$. (50)
The SO(3) generators $\Lambda^{a}~(a=1,2,3)$ are expressed as
$(\Lambda^{a})_{bc}=-i\epsilon_{abc}$ (51)
and satisfy
$[\Lambda^{a},\Lambda^{b}]=i\epsilon_{abc}\Lambda^{c},$ (52)
that is, the SU(2) algebra relation in the adjoint representation.
The general SO(3) algebra is expressed as
$\theta\cdot\Lambda\equiv\theta^{a}\Lambda^{a}\in{\rm so}(3)$ (53)
using three real parameters $\theta^{i}\in{\bf R}$.
We note the relation,
$(\hat{\theta}\cdot\Lambda)^{3}=\hat{\theta}\cdot\Lambda,$ (54)
with $\hat{\theta}^{i}\equiv\theta^{i}/\theta$ and $\theta\equiv\sqrt{\theta^{i}\theta^{i}}$. In fact, the relations
$(\Lambda^{a}\Lambda^{b})_{ik}=$ $\!\!\!\!\!\!\!\!\Lambda^{a}_{ij}\Lambda^{b}_{jk}=-\epsilon_{aij}\epsilon_{bjk}=\delta_{ab}\delta_{ik}-\delta_{ak}\delta_{bi},$ (55)
$(\Lambda^{a}\Lambda^{b}\Lambda^{c})_{il}$ $=$ $(\Lambda^{a}\Lambda^{b})_{ik}\Lambda^{c}_{kl}=-i(\delta_{ab}\delta_{ik}-\delta_{ak}\delta_{bi})\epsilon_{ckl}$ (56)
$=$ $-i(\delta_{ab}\epsilon_{cil}-\delta_{bi}\epsilon_{cal})$ (57)
lead to
$[(\hat{\theta}\cdot\Lambda)^{3}]_{il}$ $=$ $\hat{\theta}^{a}\hat{\theta}^{b}\hat{\theta}^{c}(\Lambda^{a}\Lambda^{b}\Lambda^{c})_{il}$ (58)
$=$ $-i\hat{\theta}^{a}\hat{\theta}^{b}\hat{\theta}^{c}(\delta_{ab}\epsilon_{cil}-\delta_{bi}\epsilon_{cal})$ (59)
$=$ $-i\hat{\theta}^{c}\epsilon_{cil}=\hat{\theta}^{c}\Lambda^{c}_{il}=(\hat{\theta}\cdot\Lambda)_{il}.$ (60)
Then, the SO(3) group element $u$ can be expressed as
$u$ $\equiv$ $e^{i\theta^{a}\Lambda^{a}}=e^{i\theta(\hat{\theta}\cdot\Lambda)}=\sum_{n=0}^{\infty}\frac{(i\theta)^{n}}{n!}(\hat{\theta}\cdot\Lambda)^{n}$ (61)
$=$ $1+i(\hat{\theta}\cdot\Lambda)\sin\theta+(\hat{\theta}\cdot\Lambda)^{2}(\cos\theta-1).$ (62)
In particle physics, for example, the SO(3) subgroup of SU(3) is used to describe the H dibaryon, i.e., H(uuddss), in the Skyrme model [53, 54, 55, 56] and in holographic QCD [57, 58] for the three-flavor case.
IV Factorization of SU(3) group into SO(3) and SU(3)/SO(3) ; SO(3) projection
In this section, we consider factorization of an SU(3) group into an SO(3) subgroup and an SU(3)/SO(3) coset space. Using this factorization, we also present an SO(3) projection of an SU(3) group element.
IV.1 Factorization of SU(3) group into SO(3) and SU(3)/SO(3)
In this subsection, we introduce factorization of an SU(3) group element $U$ into an SO(3) element $u$ and an SU(3)/SO(3) element $M$. Group factorization generally has some uncertainty, reflecting the non-commutative property. For instance, there are two typical types of factorization: $U=Mu$ and $U=uM$. However, once the method of factorization is determined, the procedure becomes practically unique. (Similar uncertainty appears in the factorization of the ${\rm SU}(N_{c})$ link variable into ${\rm U}(1)^{N_{c}-1}$ and ${\rm SU}(N_{c})/{\rm U}(1)^{N_{c}-1}$ variables in the maximally Abelian gauge in SU($N_{c}$) QCD [59]. Also, similar uncertainty appears in the product Ansatz used to describe the baryon-baryon interaction in the Skyrme model [60, 61].)
In this paper, we basically use factorization of an SU(3) group element $U$ into an SO(3) subgroup element $u$ and an SU(3)/SO(3) element $M$ as follow:
$U=Mu\in{\rm SU}(3)$ (63)
of which $u$ and $M$ take the definite form of
$u$ $=$ $e^{i\theta^{i}T^{i}}\in{\rm SO(3)}~~~~~~~~~~~(i=2,5,7),$ (64)
$M$ $=$ $e^{i\theta^{\bar{i}}T^{\bar{i}}}\in{\rm SU(3)/SO(3)}~~(\bar{i}=1,3,4,6,8).$ (65)
To extract $M$ from $U$, we note the useful relation
$U^{t}U=Mu\ ^{t}u^{t}M=Muu^{\dagger}M=M^{2}\in{\rm SU(3)}.$ (66)
Based on this relation, we consider the uniqueness of the factorization. In the case of $U=Mu=M^{\prime}u^{\prime}$, one finds $U^{t}U=M^{2}=M^{\prime 2}\in{\rm SU(3)}$. If $M\sim 1$, by expanding $M=e^{i\theta^{\bar{i}}T^{\bar{i}}}=1+i\theta^{\bar{i}}T^{\bar{i}}+\cdots$, $M^{2}=M^{\prime 2}$ yields $M=M^{\prime}$, which leads to $u=M^{\dagger}U=M^{\prime\dagger}U=u^{\prime}$, and therefore the factorization is found to be unique. Note also that, unlike SU(2), 1-1 is not an element of the SU(3) group.
IV.2 SO(3) projection for SU(3) group element
In this subsection, we introduce the “SO(3) projection” for an SU(3) group element $U$. For the factorization $U=Mu$, we define the SO(3) projection as the replacement of
$U=Mu\in{\rm SU(3)}~~\rightarrow~~u\in{\rm SO(3)}.$ (67)
V Maximal SO(3) gauge in SU(3) QCD
In this section, we introduce a new gauge, which we call the “maximal SO(3) gauge”, in QCD of an SU(3) gauge theory, mainly in the Euclidean metric. We divide the gluon field $A_{\mu}(x)=A_{\mu}^{a}(x)T^{a}\in{\rm su(3)}$ into the SO(3) part and the SU(3)/SO(3) part,
$A_{\mu}(x)=A_{\mu}^{\rm SO(3)}(x)+A_{\mu}^{\rm SU(3)/SO(3)}(x)\in{\rm su(3)},$~~~~~ (68)
$\left\{\begin{aligned} &A_{\mu}^{\rm SO(3)}(x)\equiv\sum_{i=2,5,7}A_{\mu}^{i}(x)T^{i}\in{\rm so}(3)\in{\rm su(3)}\cr&A_{\mu}^{\rm SU(3)/SO(3)}(x)\equiv\sum_{{\bar{i}}=1,2,4,6,8}A_{\mu}^{\bar{i}}(x)T^{\bar{i}}\in{\rm su(3)}.\end{aligned}\right$. (70)
V.1 Definition
In Euclidean SU(3) QCD, the maximal SO(3) gauge is defined to minimize
$R[A]$ $\equiv$ $\int d^{4}x\sum_{\mu}\sum_{\bar{i}=1,3,4,6,8}A_{\mu}^{\bar{i}}(x)^{2}$ (71)
$=$ $\int d^{4}x\sum_{\mu}\{A_{\mu}^{1}(x)^{2}+A_{\mu}^{3}(x)^{2}$ (73)
$~~~~~~~~~+A_{\mu}^{4}(x)^{2}+A_{\mu}^{6}(x)^{2}+A_{\mu}^{8}(x)^{2}\}$
by SU(3) gauge transformation. The functional $R[A]$ is also expressed as
$R[A]$ $=$ $4\int d^{4}x\sum_{\mu}\sum_{\bar{i}=1,3,4,6,8}[{\rm Tr}~(T^{\bar{i}}A_{\mu}(x))]^{2}$ (74)
$=$ $2\int d^{4}x\sum_{\mu}~[{\rm Tr}~A^{\rm SU(3)/SO(3)}_{\mu}(x)^{2}],$ (75)
using the “SU(3)/SO(3) part” defined by
$A_{\rm SU(3)/SO(3)}^{\mu}(x)\equiv\sum_{\bar{i}=1,3,4,6,8}A^{\mu}_{\bar{i}}(x)T^{\bar{i}}\in{\rm su(3)}.$ (76)
In fact, roughly speaking, the maximal SO(3) gauge globally minimizes the SU(3)/SO(3) part, $A_{\rm SU(3)/SO(3)}^{\mu}(x)$, in the SU(3) gauge field $A_{\mu}(x)=A_{\mu}^{a}(x)T^{a}\in{\rm su}(3)$.
V.2 Residual gauge symmetry
In the maximal SO(3) gauge, an SO(3) gauge symmetry remains, since $R[A]$ is invariant under an arbitrary SO(3) gauge transformation
$A^{\mu}\rightarrow\Omega_{\rm SO(3)}\left(A^{\mu}+\frac{1}{ig}\partial^{\mu}\right)\Omega_{\rm SO(3)}^{\dagger}$ (77)
with an SO(3) gauge function
$\Omega_{\rm SO(3)}(x)=e^{i\chi^{i}(x)T^{i}}\in{\rm SO}(3),~\chi^{i}\in{\bf R}~(i=2,5,7).$ (78)
The proof is as follows: Here, we divide the SU(3) indices $a\equiv 1,2,..,8$ into $i\equiv 2,5,7$ and $\bar{i}\equiv 1,3,4,6,8$. Since any SO(3) transformation is achieved by performing an infinite number of infinitesimal transformations, we consider infinitesimal SO(3) gauge transformation,
$\Omega_{\rm SO(3)}(x)=1+i\chi^{i}(x)T^{i},\quad|\chi^{i}(x)|\ll 1,$ (79)
up to the first order of $\chi^{i}(x)$. By this transformation, the SU(3)/SO(3) gluon component $A^{\bar{i}}_{\mu}(x)$ is transformed as
$A^{\bar{i}}_{\mu}(x)=2~{\rm Tr}~[T^{\bar{i}}A_{\mu}(x)]$ (80)
$\rightarrow$ $2~{\rm Tr}~\Big[T^{\bar{i}}~\Omega_{\rm SO(3)}\Big(A_{\mu}+\frac{1}{ig}\partial_{\mu}\Big)\Omega^{\dagger}_{\rm SO(3)}\Big]$ (81)
$=$ $2~{\rm Tr}~[T^{\bar{i}}(1+iT^{i}\chi^{i})A_{\mu}(1-iT^{j}\chi^{j})]$ (83)
$\qquad-\frac{2}{g}{\rm Tr}[T^{\bar{i}}\partial_{\mu}(T^{i}\chi^{i})]$
$=$ $A_{\mu}^{\bar{i}}(x)+2~{\rm Tr}(iT^{i}\chi^{i}[A_{\mu},T^{\bar{i}}])$ (84)
$=$ $A_{\mu}^{\bar{i}}(x)+\delta A_{\mu}^{\bar{i}}(x),$ (85)
where the term including a derivative is zero because of ${\rm Tr}(T^{\bar{i}}T^{i})=0$. Then, the variation of the SU(3)/SO(3) gluon component $A^{\bar{i}}_{\mu}(x)$ is given as
$\delta A_{\mu}^{\bar{i}}(x)$ $\equiv$ $2{\rm Tr}(iT^{i}\chi^{i}[A_{\mu},T^{\bar{i}}])=2i\chi^{i}A_{\mu}^{a}{\rm Tr}(T^{i}[T^{a},T^{\bar{i}}])$ (86)
$=$ $-2\chi^{i}A_{\mu}^{a}f^{a\bar{i}b}{\rm Tr}(T^{i}T^{b})=f^{i\bar{i}a}\chi^{i}A_{\mu}^{a}$ (87)
$=$ $f^{i\bar{i}j}\chi^{i}A_{\mu}^{j}+f^{i\bar{i}\bar{j}}\chi^{i}A_{\mu}^{\bar{j}}=f^{i\bar{i}\bar{j}}\chi^{i}A_{\mu}^{\bar{j}},$ (88)
because of $f^{\bar{i}ij}=0$, as was shown in Eq.(40). For the infinitesimal SO(3) gauge transformation, the variation of $R[A]$ is written as
$\delta R[A]$ $\equiv$ $R[A+\delta A]-R[A]=2\int d^{4}x~A_{\mu}^{\bar{i}}(x)~\delta A_{\mu}^{\bar{i}}(x)$ (89)
$=$ $2\int d^{4}x~A_{\mu}^{\bar{i}}(x)\cdot f^{i\bar{i}\bar{j}}\chi^{i}(x)A_{\mu}^{\bar{j}}(x)$ (90)
$=$ $2\int d^{4}x~\chi^{i}(x)f^{i\bar{i}\bar{j}}A_{\mu}^{\bar{i}}(x)A_{\mu}^{\bar{j}}(x)=0.$ (91)
Since any SO(3) transformation is achieved through an infinite number of infinitesimal transformations, $R[A]$ is unchanged by any SO(3) gauge transformation.
Thus, the maximal SO(3) gauge is a type of partial gauge fixing, and QCD in this gauge behaves as an SO(3) gauge theory. (This is similar to the maximally Abelian gauge, which preserves Abelian gauge symmetry. QCD in this gauge behaves as an Abelian gauge theory [59, 62, 63, 64, 65, 66, 67].)
V.3 Local gauge-fixing condition
Using the SO(3) covariant derivative $D_{\rm SO(3)}^{\mu}$ defined by
$D_{\rm SO(3)}^{\mu}\equiv\partial^{\mu}+igA_{\rm SO(3)}^{\mu}=\partial^{\mu}+ig\sum_{i=2,5,7}A_{i}^{\mu}T^{i},$ (92)
the local gauge-fixing condition of the maximal SO(3) gauge is expressed as
$\left[D^{\mu}_{\rm SO(3)},A_{\mu}^{\rm SU(3)/SO(3)}\right]=0,$ (93)
or equivalently,
$\Big[\partial^{\mu}+ig\sum_{i=2,5,7}A^{\mu}_{i}T^{i},\sum_{{\bar{i}}=1,3,4,6,8}A_{\mu}^{\bar{i}}T^{\bar{i}}\Big]=0.$ (94)
Here, $A_{\rm SO(3)}^{\mu}(x)$ and $A^{\mu}_{\rm SU(3)/SO(3)}(x)$ are the SO(3) part $A_{\rm SO(3)}^{\mu}(x)\equiv\sum_{i=2,5,7}A^{\mu}_{i}(x)T^{i}$ and the SU(3)/SO(3) part $A_{\rm SU(3)/SO(3)}^{\mu}(x)\equiv\sum_{\bar{i}=1,3,4,6,8}A^{\mu}_{\bar{i}}(x)T^{\bar{i}}$ of the gluon field $A^{\mu}(x)\in{\rm su(3)}$, respectively. According to the SO(3) covariant derivative, the SO(3) gauge symmetry is found to remain in this local gauge-fixing condition.
Using the SO(3) index $i$ and SU(3)/SO(3) index $\bar{i}$}, the local gauge-fixing condition (94) is expressed as
$\partial^{\mu}A_{\mu}^{\bar{i}}+gf^{i\bar{i}\bar{j}}A^{\mu}_{i}A_{\mu}^{\bar{j}}=0.$ (95)
Here, we have used the algebra relation,
$[T^{i},T^{\bar{i}}]=if^{i\bar{i}a}T^{a}=if^{i\bar{i}j}T^{j}+if^{i\bar{i}\bar{j}}T^{\bar{j}}=if^{i\bar{i}\bar{j}}T^{\bar{j}},$ (96)
due to $f^{\bar{i}ij}=0$ in Eq. (40).
The derivation of the local gauge-fixing condition (93) is as follows. As was shown in the previous subsection, the maximal SO(3) gauge is SO(3) gauge invariant; therefore, it is sufficient to consider the SU(3)/SO(3) gauge transformation. We consider infinitesimal SU(3)/SO(3) gauge transformation,
$\Omega_{\rm SU(3)/SO(3)}(x)$ $=$ $e^{i\chi^{\bar{i}}(x)T^{\bar{i}}}$ (97)
$\simeq$ $1+i\chi^{\bar{i}}(x)T^{\bar{i}},\quad|\chi^{\bar{i}}(x)|\ll 1,$~~~ (98)
up to the first order of $\chi^{\bar{i}}(x)$. By this transformation, the SU(3)/SO(3) gluon component $A^{\bar{i}}_{\mu}(x)$ is transformed as
$A^{\bar{i}}_{\mu}(x)=2~{\rm Tr}~[T^{\bar{i}}A_{\mu}(x)]$ (99)
$\rightarrow$ $2~{\rm Tr}~\Big[T^{\bar{i}}~\Omega_{\rm SU(3)/SO(3)}\Big(A_{\mu}+\frac{1}{ig}\partial_{\mu}\Big)\Omega^{\dagger}_{\rm SU(3)/SO(3)}\Big]$ (100)
$=$ $2~{\rm Tr}~[T^{\bar{i}}(1+iT^{\bar{j}}\chi^{\bar{j}})A_{\mu}(1-iT^{\bar{k}}\chi^{\bar{k}})]$ (102)
$\qquad-\frac{2}{g}{\rm Tr}[T^{\bar{i}}\partial_{\mu}(T^{\bar{k}}\chi^{\bar{k}})]$
$=$ $A_{\mu}^{\bar{i}}(x)+2{\rm Tr}(iT^{\bar{j}}\chi^{\bar{j}}[A_{\mu},T^{\bar{i}}])-\frac{1}{g}\partial_{\mu}\chi^{\bar{i}}$ (103)
$=$ $A_{\mu}^{\bar{i}}(x)+\delta A_{\mu}^{\bar{i}}(x).$ (104)
Then, the variation of the SU(3)/SO(3) gluon component $A_{\mu}^{\bar{i}}(x)$ is expressed as
$\delta A_{\mu}^{\bar{i}}(x)$ $\equiv$ $2{\rm Tr}(iT^{\bar{j}}\chi^{\bar{j}}[A_{\mu},T^{\bar{i}}])-\frac{1}{g}\partial_{\mu}\chi^{\bar{i}}$ (105)
$=$ $2i\chi^{\bar{j}}A_{\mu}^{a}{\rm Tr}(T^{\bar{j}}[T^{a},T^{\bar{i}}])-\frac{1}{g}\partial_{\mu}\chi^{\bar{i}}$ (106)
$=$ $-\chi^{\bar{j}}A_{\mu}^{a}f^{a\bar{i}\bar{j}}-\frac{1}{g}\partial_{\mu}\chi^{\bar{i}}$ (107)
$=$ $-\chi^{\bar{j}}A_{\mu}^{i}f^{i\bar{i}\bar{j}}-\frac{1}{g}\partial_{\mu}\chi^{\bar{i}}$ (108)
because of $f^{\bar{i}\bar{j}\bar{k}}=0$ in Eq.(41). Then, one finds
$A_{\mu}^{\bar{i}}(x)\delta A_{\mu}^{\bar{i}}(x)$ $=$ $-A_{\mu}^{\bar{i}}(\frac{1}{g}\partial_{\mu}\chi^{\bar{i}}+\chi^{\bar{j}}A_{\mu}^{i}f^{i\bar{i}\bar{j}})$ (109)
$=$ $-\frac{1}{g}\partial_{\mu}\chi^{\bar{i}}\cdot A_{\mu}^{\bar{i}}+\chi^{\bar{i}}f^{i\bar{i}\bar{j}}A_{\mu}^{i}A_{\mu}^{\bar{j}},$~~~~ (110)
where the indices $\bar{i}$} and $\bar{j}$} are interchanged in part. For the infinitesimal SU(3) gauge transformation, the variation of $R[A]$ is written as
$\delta R[A]\equiv R[A+\delta A]-R[A]=2\int d^{4}x~A_{\mu}^{\bar{i}}(x)~\delta A_{\mu}^{\bar{i}}(x)$ (111)
$=2\int d^{4}x~\Big(-\frac{1}{g}\partial_{\mu}\chi^{\bar{i}}\cdot A_{\mu}^{\bar{i}}+\chi^{\bar{i}}f^{i\bar{i}\bar{j}}A_{\mu}^{i}A_{\mu}^{\bar{j}}\Big)$ (112)
$=\frac{2}{g}\int d^{4}x~\chi^{\bar{i}}(x)(\partial_{\mu}A_{\mu}^{\bar{i}}+gf^{i\bar{i}\bar{j}}A_{\mu}^{i}A_{\mu}^{\bar{j}}),$ (113)
where a partial integration is performed with the boundary condition that the gauge function vanishes at infinity, $\chi^{\bar{i}}(|x|=\infty)=0$. The extreme condition of $R[A]$, $\delta R=0$, with respect to the infinitesimal gauge transformation with an arbitrary small $\chi^{\bar{i}}(x)$ leads to $\partial_{\mu}A_{\mu}^{\bar{i}}+gf^{i\bar{i}\bar{j}}A_{\mu}^{i}A_{\mu}^{\bar{j}}=0$, that is, the local gauge-fixing condition (95).
VI Maximal SO(3) gauge in SU(3) lattice QCD
In this section, we formulate the “maximal SO(3) gauge” in SU(3) lattice QCD.
VI.1 Definition
For the SU(3) link-variable $U_{\mu}(s)$, we consider the factorization of
$U_{\mu}(s)=M_{\mu}(s)u_{\mu}(s)\in{\rm SU(3)}$ (114)
in terms of SO(3) and SU(3)/SO(3) variables,
$u_{\mu}(s)$ $\equiv$ $e^{iT^{i}\theta_{\mu}^{i}(s)}\in{\rm SO(3)},$ (115)
$M_{\mu}(s)$ $\equiv$ $e^{iT^{\bar{i}}\theta_{\mu}^{\bar{i}}(s)}\in{\rm SU(3)/SO(3)}.$ (116)
Due to $^{t}u_{\mu}(s)=u_{\mu}^{\dagger}(s)$ and $^{t}M_{\mu}(s)=M_{\mu}(s)$, we find
$U_{\mu}(s)^{t}U_{\mu}(s)=M_{\mu}(s)^{2},$ (117)
which is the lattice QCD version of Eq. (66).
In Euclidean SU(3) lattice QCD, the maximal SO(3) gauge is defined to maximize
$R_{L}[U]$ $\equiv$ $\sum_{s,\mu}{\rm Re}~{\rm Tr}\{U_{\mu}(s)^{t}U_{\mu}(s)\}$ (118)
$=$ $\sum_{s,\mu}{\rm Re}~{\rm Tr}\{M_{\mu}(s)^{2}\}$ (119)
by SU(3) gauge transformation.
Thus, in the maximal SO(3) gauge, the SU(3)/SO(3) factor $M_{\mu}(s)$ in the link-variable $U_{\mu}(s)\in{\rm SU(3)}$ is maximally close to unity, since 1-1 is not an SU(3) group element. Therefore, one roughly expects
$M_{\mu}(s)\sim 1,~~{\rm i.e.,}~~U_{\mu}(s)\sim u_{\mu}(s)$ (120)
in the maximal SO(3) gauge.
VI.2 Continuum limit
We examine the continuum limit of the lattice QCD definition of the maximal SO(3) gauge. In the maximal SO(3) gauge, $M_{\mu}(s)=e^{iT^{\bar{i}}\theta_{\mu}^{\bar{i}}(s)}\in{\rm SU(3)/SO(3)}$ is close to unity, and $\theta_{\mu}^{\bar{i}}(s)$ is considered to be small, which allows the following expansion:
$U_{\mu}(s)^{t}U_{\mu}(s)$ $=$ $M_{\mu}(s)^{2}=e^{2iT^{\bar{i}}\theta_{\mu}^{\bar{i}}(s)}$ (121)
$\simeq$ $1+2iT^{\bar{i}}\theta_{\mu}^{\bar{i}}(s)-2T^{\bar{i}}\theta_{\mu}^{\bar{i}}(s)T^{\bar{j}}\theta_{\mu}^{\bar{j}}(s),$~~~~ (122)
and then
$R_{L}[U]$ $\equiv$ $\sum_{s,\mu}{\rm Re~Tr}~\{U_{\mu}(s)^{t}U_{\mu}(s)\}$ (123)
$\simeq$ $-\sum_{s,\mu}{\rm Re~Tr}~\theta_{\mu}^{\bar{i}}(s)\theta_{\mu}^{\bar{i}}(s)+{\rm const.}$ (124)
Therefore, in the continuum limit, maximizing $R_{L}[U]$ corresponds to minimizing $R[A]$ defined in Eq.(73). This yields the definition of the maximal SO(3) gauge in continuum SU(3) QCD, presented in Sec. V.
VI.3 Residual gauge symmetry
In the maximal SO(3) gauge, an SO(3) gauge symmetry remains, since $R_{L}[U]$ is invariant under an arbitrary SO(3) gauge transformation
[…]
with an SO(3) gauge function $v(s)\in{\rm SO(3)}$. In fact, by the SO(3) gauge transformation, $R_{L}[U]$ is unchanged as
[…]
$~~~~~~~~~~~\cdot^{t}v^{\dagger}(s+\mu)^{t}U_{\mu}(s)^{t}v(s)\}$
$=$ $\sum_{s,\mu}{\rm Re}~{\rm Tr}\{U_{\mu}(s)^{t}U_{\mu}(s)\}=R_{L}[U],$ (128)
because of the SO(3) property, $^{t}v(s)=v^{\dagger}(s)$.
For $U_{\mu}(s)=M_{\mu}(s)u_{\mu}(s)$, we consider the SO(3) gauge transformation of $M_{\mu}(s)\in{\rm SU(3)/SO(3)}$ and $u_{\mu}(s)\in{\rm SO(3)}$. To keep the factorial form, we find
$M_{\mu}(s)$ $\rightarrow$ $M_{\mu}^{v}(s)=v(s)M_{\mu}(s)v^{\dagger}(s)\in{\rm SU(3)/SO(3)},$ (129)
$u_{\mu}(s)$ $\rightarrow$ $u_{\mu}^{v}(s)=v(s)u_{\mu}(s)v^{\dagger}(s+\hat{\mu})\in{\rm SO(3)}.$ (130)
In fact, the SU(3)/SO(3) property of $^{t}M(s)=M(s)$ is kept by the SO(3) gauge transformation as
[…]
$=$ $v(s)M_{\mu}(s)v^{\dagger}(s)=M^{v}_{\mu}(s)$ (132)
because of $^{t}v(s)=v(s)^{\dagger}\in{\rm SO(3)}$.
Thus, by the SO(3) gauge transformation, $M_{\mu}(s)\in{\rm SU(3)/SO(3)}$ obeys a unitary transformation at the same spacetime, and $u_{\mu}(s)\in{\rm SO(3)}$ behaves as a gauge variable that links neighboring spacetime sites. (This situation is similar to the behavior of diagonal and off-diagonal gluons by residual Abelian gauge transformation in the maximally Abelian gauge [59, 65, 66, 67].)
VI.4 SU(3) gauge transformation property
Now, we consider general SU(3) gauge transformation of the quantity $R_{L}[U]$ defined in Eq.(119). We factorize the gauge function $V(s)\in{\rm SU(3)}$ as
$V(s)$ $=$ $v(s)\bar{V}(s)\quad{\rm or}\quad V^{\dagger}(s)=\bar{V}^{\dagger}(s)v^{\dagger}(s),$ (133)
$v(s)$ $\in$ ${\rm SO(3)},\quad\bar{V}(s)\in{\rm SU(3)/SO(3)},$ (134)
which satisfy $^{t}v_{\mu}(s)=v_{\mu}^{\dagger}(s)$ and $^{t}\bar{V}_{\mu}(s)=\bar{V}_{\mu}(s)$.
Then, the SU(3) link variable $U_{\mu}(s)$ is gauge transformed as
[…]
and then $R_{L}[U]$ is transformed as
$R_{L}[U]\equiv\sum_{s,\mu}{\rm Re}~{\rm Tr}\{U_{\mu}(s)^{t}U_{\mu}(s)\}$ (137)
$\rightarrow$ $\sum_{s,\mu}{\rm Re}~{\rm Tr}\{\bar{V}(s)^{2}~U_{\mu}(s)~\bar{V}^{\dagger}(s+\hat{\mu})^{2}\ {}^{t}U_{\mu}(s)\}.$~~~~~ (138)
Since the gauge transformation of $R_{L}[U]$ is independent of the SO(3) gauge function $v(s)$, we find again that the SO(3) part remains as a residual gauge symmetry.
VI.5 Local SU(3) gauge transformation
For $R_{L}[U]$ defined in Eq. (119), we examine the local SU(3) gauge transformation consisting of a nontrivial transformation only at one site $s$ in lattice QCD,
$V(s^{\prime})=V\delta_{s^{\prime}s},\quad V\in{\rm SU(3)}.$ (139)
As a result of this gauge transformation, $R_{L}[U]$ differs by $\Delta R_{L}$:
$\Delta R_{L}\equiv R_{L}[U^{V}]-R_{L}[U]$ (140)
$=$ $\sum_{\mu=1}^{4}{\rm ReTr}\{VU_{\mu}(s)^{t}U_{\mu}(s)^{t}V\}$ (141)
[…]
$=$ ${\rm ReTr}\{^{t}VV\sum_{\mu=1}^{4}\sum_{\pm}U_{\pm\mu}(s)^{t}U_{\pm\mu}(s)\}.$ (143)
Here, we have used $U_{-\mu}(s)\equiv U_{\mu}(s-\mu)^{\dagger}$ and ${\rm ReTr}(M)={\rm ReTr}(M^{\dagger})$ for any 3 $\times$ 3 matrix $M$.
Applying the factorization of $V=v\bar{V}$ (or $V^{\dagger}=\bar{V}^{\dagger}v^{\dagger}$) for the gauge function, we find
$^{t}VV$ $=$ $\bar{V}v^{\dagger}v\bar{V}=\bar{V}^{2}$ (144)
and obtain
$\Delta R_{L}$ $=$ ${\rm ReTr}\{{\bar{V}}^{2}\sum_{\mu=1}^{4}\sum_{\pm}U_{\pm\mu}(s)^{t}U_{\pm\mu}(s)\}.$ (145)
This is useful for the practical gauge-fixing procedure of the maximal SO(3) gauge in SU(3) lattice QCD.
VII SO(3) real algebra method in SU(3) QCD
In this section, we present the formalism of the “SO(3) real algebra method” for continuum SU(3) QCD at finite baryon-number densities.
Our strategy is to divide the SO(3) part, which is essentially real, from the SU(3) fermionic determinant in the maximal SO(3) gauge. The QCD generating functional in the maximal SO(3) gauge at finite densities is expressed as
$Z_{\rm SO(3)}(\mu)$ $=$ $\int DAe^{-S_{G}[A]}~\delta^{\rm GF}_{\rm SO(3)}[A]~\Delta_{\rm SO(3)}^{\rm FP}[A]~{\rm Det}(\not\!\!D+m+\mu\gamma_{4})$ (146)
$=$ $\int DAe^{-S_{G}[A]}~\delta^{\rm GF}_{\rm SO(3)}[A]~\Delta_{\rm SO(3)}^{\rm FP}[A]~{\rm Det}(\not\!\!D_{\rm SO(3)}+m+\mu\gamma_{4})\cdot\frac{{\rm Det}(\not\!\!D+m+\mu\gamma_{4})}{{\rm Det}(\not\!\!D_{\rm SO(3)}+m+\mu\gamma_{4})},$ (147)
where $\delta^{\rm GF}_{\rm SO(3)}[A]$ denotes the maximal SO(3) gauge-fixing constraint and $\Delta_{\rm SO(3)}^{\rm FP}[A]$ is the corresponding Faddeev-Popov determinant. $D_{\rm SO(3)}^{\mu}$ denotes the SO(3) covariant derivative defined in Eq.(92), $D_{\rm SO(3)}^{\mu}\equiv\partial^{\mu}+igA_{\rm SO(3)}^{\mu}$.
In Eq. (147), the following factor $p[A]$ is generally real, and $p[A]$ is non-negative for the even-number flavor case ($N_{f}=2n$) of the same quark mass:
$p[A]\equiv e^{-S_{G}[A]}~\delta^{\rm GF}_{\rm SO(3)}[A]~\Delta_{\rm SO(3)}^{\rm FP}[A]~{\rm Det}(\not\!\!D_{\rm SO(3)}+m+\mu\gamma_{4})$ $\in$ ${\bf R}$ (148)
$\geq$ $0\quad(N_{f}=2n).$ (149)
Therefore, for the $N_{f}=2n$ case, this non-negative factor $p[A]$ can be identified as a probability factor in the Monte Carlo method.
Using the relation
$D^{\mu}=D^{\mu}_{\rm SO(3)}+igA^{\mu}_{\rm SU(3)/SO(3)}=D^{\mu}_{\rm SO(3)}+igA^{\mu}_{\bar{i}}T^{\bar{i}},$~~~ (150)
the SU(3) fermionic determinant is expressed as
${\rm ln~Det}(\not\!\!D+m+\mu\gamma_{4})$ (151)
$=$ ${\rm ln~Det}(\not\!\!D_{\rm SO(3)}+m+\mu\gamma_{4}+ig\not\!\!A_{\rm SU(3)/SO(3)})$ (152)
$=$ ${\rm Tr~ln}(\not\!\!D_{\rm SO(3)}+m+\mu\gamma_{4}+ig\not\!\!A_{\rm SU(3)/SO(3)})$ (153)
$=$ ${\rm Tr~ln}(\not\!\!D_{\rm SO(3)}+m+\mu\gamma_{4})$ (154)
$+$ ${\rm Tr~ln}\left[1+ig\not\!\!A_{\rm SU(3)/SO(3)}\frac{1}{\not\!\!D_{\rm SO(3)}+m+\mu\gamma_{4}}\right].$~~~~~ (155)
Then, the fermionic determinant ratio in Eq.(147) can be written as
${\rm ln~}\frac{{\rm Det}(\not\!\!D+m+\mu\gamma_{4})}{{\rm Det}(\not\!\!D_{\rm SO(3)}+m+\mu\gamma_{4})}={\rm ln~Det}(\not\!\!D+m+\mu\gamma_{4})-{\rm ln~Det}(\not\!\!D_{\rm SO(3)}+m+\mu\gamma_{4})$ (156)
$=$ ${\rm Tr~ln}\left[1+ig\not\!\!A_{\rm SU(3)/SO(3)}\frac{1}{\not\!\!D_{\rm SO(3)}+m+\mu\gamma_{4}}\right]=-\sum_{n=1}^{\infty}\frac{(-ig)^{n}}{n}{\rm Tr}\left(\not\!\!A_{\rm SU(3)/SO(3)}\frac{1}{\not\!\!D_{\rm SO(3)}+m+\mu\gamma_{4}}\right)^{n}$ (157)
$=$ $\sum_{n=1}^{\infty}\frac{(-)^{n+1}g^{2n}}{2n}{\rm Tr}\left(\not\!\!A_{\rm SU(3)/SO(3)}\frac{1}{\not\!\!D_{\rm SO(3)}+m+\mu\gamma_{4}}\right)^{2n}+i\sum_{n=1}^{\infty}\frac{(-)^{n}g^{2n+1}}{2n+1}{\rm Tr}\left(\not\!\!A_{\rm SU(3)/SO(3)}\frac{1}{\not\!\!D_{\rm SO(3)}+m+\mu\gamma_{4}}\right)^{2n+1}.$ (158)
This series consists of the SU(3)/SO(3) gluon, $A_{\rm SU(3)/SO(3)}^{\mu}$, and the quark propagator interacting with the SO(3) gluon, $A_{\rm SO(3)}^{\mu}$. Note that the SU(3)/SO(3) gluon $A_{\rm SU(3)/SO(3)}^{\mu}$ is real,
$A_{\rm SU(3)/SO(3)}^{\mu}(x)^{*}=A_{\rm SU(3)/SO(3)}^{\mu}(x),$ (160)
and the SO(3) covariant derivative $D_{\rm SO(3)}^{\mu}$ is also real,
${D_{\rm SO(3)}^{\mu}}^{*}=(\partial^{\mu}+igA_{\rm SO(3)}^{\mu})^{*}=D_{\rm SO(3)}^{\mu},$ (161)
because of ${A_{\rm SO(3)}^{\mu}}^{*}=-A_{\rm SO(3)}^{\mu}$. Then, we find for real $\mu$
$\left[{\rm Tr}\left(\not\!\!A_{\rm SU(3)/SO(3)}\frac{1}{\not\!\!D_{\rm SO(3)}+m+\mu\gamma_{4}}\right)^{n}\right]^{*}$ (162)
$=$ ${\rm Tr}\left(C\not\!\!A_{\rm SU(3)/SO(3)}\frac{1}{\not\!\!D_{\rm SO(3)}+m+\mu\gamma_{4}}C^{-1}\right)^{n}$ (163)
$=$ ${\rm Tr}\left(\not\!\!A_{\rm SU(3)/SO(3)}\frac{1}{\not\!\!D_{\rm SO(3)}+m+\mu\gamma_{4}}\right)^{n}\in{\bf R},$ (164)
where $C$ is the complex conjugation matrix satisfying $\gamma_{\mu}^{*}=C\gamma_{\mu}C^{-1}$, i.e., Eq. (16). Therefore, the fermionic determinant ratio is expressed as
${\rm ln~}\frac{{\rm Det}(\not\!\!D+m+\mu\gamma_{4})}{{\rm Det}(\not\!\!D_{\rm SO(3)}+m+\mu\gamma_{4})}$ $=$ $p+iq,$ (165)
${\rm i.e.,}\quad\frac{{\rm Det}(\not\!\!D+m+\mu\gamma_{4})}{{\rm Det}(\not\!\!D_{\rm SO(3)}+m+\mu\gamma_{4})}$ $=$ $e^{p+iq},$ (166)
where $p$ and $q$ are the real numbers defined by
$p\equiv\sum_{n=1}^{\infty}\frac{(-)^{n+1}g^{2n}}{2n}{\rm Tr}\left(\not\!\!A_{\rm SU(3)/SO(3)}\frac{1}{\not\!\!D_{\rm SO(3)}+m+\mu\gamma_{4}}\right)^{2n}\in{\bf R},$ (167)
$q\equiv\sum_{n=1}^{\infty}\frac{(-)^{n}g^{2n+1}}{2n+1}{\rm Tr}\left(\not\!\!A_{\rm SU(3)/SO(3)}\frac{1}{\not\!\!D_{\rm SO(3)}+m+\mu\gamma_{4}}\right)^{2n+1}\in{\bf R}.$ (168)
For the fermionic determinant ratio, the real number $p$ yields a harmless real factor of $e^{p}$, and $q$ gives the phase factor $e^{iq}$, which has the potential to cause the sign problem in the Monte Carlo method.
If the norm of $A^{\mu}_{\rm SU(3)/SO(3)}$ is well suppressed in the maximally SO(3) gauge, $q$ may not fluctuate significantly over the sampled gauge configurations, and also the infinite series in Eq. (LABEL:eq:expansion) could be truncated to a finite one. Even after the truncation, the effect of the chemical potential $\mu$ is firmly included in the SO(3) fermionic determinant.
VIII SO(3) real algebra method
in SU(3) lattice QCD
Now, we formulate the SO(3) real algebra method in SU(3) lattice QCD at a finite density with a quark-number chemical potential $\mu$. Using the SU(3) lattice fermion kernel $K[U;\mu]$, the SU(3) fermionic determinant is expressed as ${\rm Det}~K[U;\mu]$. The lattice QCD generating functional in the maximal SO(3) gauge at finite densities is expressed as
$Z_{\rm SO(3)}(\mu)$ $=$ $\int DUe^{-S_{G}[U]}~\delta^{\rm GF}_{\rm SO(3)}[U]~\Delta_{\rm SO(3)}^{\rm FP}[U]~{\rm Det}~K[U;\mu]$ (169)
$=$ $\int DUe^{-S_{G}[U]}~\delta^{\rm GF}_{\rm SO(3)}[U]~\Delta_{\rm SO(3)}^{\rm FP}[U]~{\rm Det}~K[u;\mu]\cdot\frac{{\rm Det}~K[U;\mu]}{{\rm Det}~K[u;\mu]},$ (170)
where $\delta^{\rm GF}_{\rm SO(3)}[U]$ denotes the maximal SO(3) gauge-fixing constraint and $\Delta_{\rm SO(3)}^{\rm FP}[U]$ is the corresponding Faddeev-Popov determinant in lattice QCD.
In Eq. (170), the following factor $p[U]$ is generally real, and $p[U]$ is non-negative for the even-number flavor case ($N_{f}=2n$) of the same quark mass:
$p[U]\equiv e^{-S_{G}[U]}~\delta^{\rm GF}_{\rm SO(3)}[U]~\Delta_{\rm SO(3)}^{\rm FP}[U]~{\rm Det}~K[u;\mu]$ $\in$ ${\bf R}$ (171)
$\geq$ $0\quad(N_{f}=2n).$ (172)
Therefore, for the $N_{f}=2n$ case, the non-negative factor $p[U]$ can be identified as a probability factor.
The SO(3) real algebra method employs the non-negative factor $p[U]$ as a probability factor of Monte Carlo calculations for the $N_{f}=2n$ case with the same quark mass. In this method, SU(3) gauge configurations $\{U_{k}\}_{k=1,2,...,N}$ can be generated by the most importance sampling with the weight of $p[U]$. This procedure is achieved through the Monte Carlo update for the SU(3) action factor $e^{-S_{G}[U]}$ and the SO(3) fermionic determinant ${\rm Det}~K[u;\mu]$, while applying the maximal SO(3) gauge fixing. In fact, the SO(3) real algebra method alternates between Monte Carlo updates and the maximal SO(3) gauge fixing.
For an arbitrary operator $O[U]$, its expectation value is evaluated as
$\langle O[U]\rangle_{\rm SU(3)}$ $=$ $\left\langle O[U]\cdot\frac{{\rm Det}~K[U;\mu]}{{\rm Det}~K[u;\mu]}\right\rangle_{\rm SO(3)}$ (173)
$=$ $\sum_{k=1}^{N}O[U_{k}]\cdot\frac{{\rm Det}~K[U_{k};\mu]}{{\rm Det}~K[u_{k};\mu]},$ (174)
where $U_{k}=M_{k}u_{k}$ denotes the $k$ th gauge configuration. Here, $\langle...\rangle_{\rm SO(3)}$ means the expectation value with the weight $p[U]$, which is real and non-negative for the case of $N_{f}=2n$ with the same quark mass. In fact, the SO(3) real algebra method is a type of reweighting method [26, 27, 28]. Note that this method only uses gauge fixing during the most importance sampling, and the final result is completely gauge invariant.
The SO(3) real algebra method can be easily extended to finite temperature QCD by imposing temporal (anti)periodicity [13]. In fact, this method can be used for SU(3) lattice QCD at finite densities and temperatures.
For arbitrary hermite operators, the expectation value $\langle O[U]\rangle_{\rm SU(3)}$ itself takes a real physical value. The problem is that there could be large fluctuations in the phase factor of the determinant ratio, which would cause numerical Monte Carlo simulations to fail.
However, if one can perform good sampling and suppress the fluctuation of the phase factor in each sampled configuration, the Monte Carlo calculation will be possible. Although the SU(3) fermionic determinant ${\rm Det}~K[U_{k};\mu]$ is gauge invariant, its numerical evaluation depends on the sampled gauge configurations, and there are possibilities for improvement in the sampling process.
In fact, if the phase factor of the determinant ratio ${\rm Det}~K[U_{k};\mu]/{\rm Det}~K[u_{k};\mu]$ is not significantly fluctuating in the Monte Carlo sampled gauge configurations, then lattice QCD calculations at finite densities would be feasible for the $N_{f}=2n$ case.
On this point, in the maximal SO(3) gauge, the SU(3) gauge configurations are expected to be maximally close to the SO(3) ones, and hence the phase factor of the determinant ratio ${\rm Det}~K[U_{k};\mu]/{\rm Det}~K[u_{k};\mu]$ in the sampled gauge configurations may not fluctuate significantly during the most importance sampling in the Monte Carlo method.
The next important step is to check the phase factor of the determinant ratio ${\rm Det}~K[U_{k};\mu]/{\rm Det}~K[u_{k};\mu]$ in the maximal SO(3) gauge in an actual lattice QCD Monte Carlo simulation. The degree of fluctuation would depend on the lattice volume, the chemical potential $\mu$, the fermion mass $m$ and so on. If there is a window that is close to the physical situation, this approach would be realistic for exploring finite density QCD.
IX Summary and conclusion
In this paper, we have proposed the SO(3) real algebra method for lattice QCD calculations at finite baryon-number densities. In this method, we have divided the SU(3) gauge variable into the SO(3) and SU(3)/SO(3) parts. Then, we have introduced the maximal SO(3) gauge by minimizing the SU(3)/SO(3) part of the SU(3) gauge variable. We have considered the SO(3) fermionic determinant, that is, the fermionic determinant of the SO(3) part of the SU(3) gauge variable.
We have noted that the SO(3) fermionic determinant is real, and it is non-negative for the even-number flavor case ($N_{f}=2n$) of the same quark mass, e.g., $m_{u}=m_{d}$. In the Monte Carlo calculation, as well as the positive SU(3) gauge action factor $e^{-S_{G}}$, the SO(3) real algebra method has used the SO(3) fermionic determinant in the maximal SO(3) gauge for the $N_{f}=2n$ case.
The SO(3) real algebra method alternates between the maximal SO(3) gauge fixing and Monte Carlo updates on the SO(3) determinant and $e^{-S_{G}}$. After the most importance sampling, we have treated the remaining ratio of the SU(3) and SO(3) fermionic determinants as a weight factor in the reweighting method. In fact, we have only used gauge fixing for the most importance sampling, and the final result is fully gauge invariant. If the phase factor of the ratio does not fluctuate significantly in the sampled gauge configurations, then lattice QCD calculations at finite densities would be feasible. This method can be extended to finite temperature QCD by imposing temporal (anti)periodicity. In fact, this method can be used to study QCD at finite densities and temperatures.
We summarize the lattice QCD procedure of the SO(3) real algebra method for finite density QCD:
  1. 1.
    Perform maximal SO(3) gauge fixing to maximize $\sum_{\mu=1}^{4}{\rm ReTr}\{U_{\mu}(s)^{t}U_{\mu}(s)\}$.
  2. 2.
    Perform a Monte Carlo calculation for the SU(3) action factor $e^{-S_{G}[U]}$ and the SO(3) fermionic determinant ${\rm Det}~K[u;\mu]$.
  3. 3.
    Repeat steps 1 and 2 alternately until thermalization is achieved. Then, sample the SU(3) gauge configurations $\{U_{k}\}_{k=1,2,...,N}$.
  4. 4.
    Use the sampled SU(3) gauge configurations in the above and include the fermionic determinant ratio ${\rm Det}~K[U;\mu]/{\rm Det}~K[u;\mu]$ as a weight factor when calculating the expectation value.
The remaining serious problem is the possibility of significant fluctuations in the phase factor of the fermionic determinant ratio, although the SU(3) gauge variable is expected to closely resemble its SO(3) part in the maximal SO(3) gauge. As a next important work, we will investigate the the phase factor distribution of the fermionic determinant ratio in the Monte Carlo sampled configurations using the SO(3) real algebra method. The degree of fluctuation would depend on the lattice volume, the chemical potential $\mu$, the temperature $T$, the fermion mass $m$ and so on.
If there is a window that is close to the physical situation, this approach would be realistic for lattice calculations of QCD at finite densities and temperatures, and the QCD phase diagram would be obtained.
The series in Eq. (LABEL:eq:expansion) may be useful. Note again that this series consists of the SU(3)/SO(3) gluon field and the quark propagator interacting with the SO(3) gluon, which are both essentially real in the trace.
Compared to other methods, the SO(3) real algebra method is a type of reweighting method [26, 27, 28]. This method is philosophically similar to the Lefschetz thimble method [33, 34, 35, 36, 37, 38], which reduces phase fluctuations by taking a suitable complex path, while this method reduces them by taking a suitable gauge. Unlike the Taylor expansion method [18, 19], this method does not use an expansion by $\mu$, so its applicable region is not limited to small $\mu$. Even if this method alone may not sufficiently suppress phase fluctuations, combining it with the other methods mentioned in Sec. I may enable the numerical simulation of finite density QCD feasible.
Finally, we briefly provide some related perspectives on the maximal SO(3) real algebra method and the maximal SO(3) gauge. First, the generalization of this method to SU($N$) gauge theories is straightforward by performing a similar procedure that divides the SU($N$) gauge variable into SO($N$) and SU(NN)/SO(NN) parts. Second, it is interesting to investigate “SO(3) dominance” in the maximal SO(3) gauge, motivated by Abelian dominance [59, 64, 65, 66, 67, 68, 69, 70] in the maximally Abelian gauge.
Acknowledgements
We thank Dr. T. M. Doi for his useful advice on the direct calculation of the fermionic determinant. K.T. is supported by JSPS Research Fellowship for Young Scientists.